# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v1

@@ 4.3 Ablation Study
$62.7$
$72.3$
$72.8$
$82.7$
$64.3$
$73.5$
$73.6$
$83.4$
$66.2$
$73.9$
$76.4$
$84.6$
$67.0$
$74.1$
$77.2$
$84.8$
$67.6$
$74.2$
$77.7$
$\mathbf{85.2}$
$\mathbf{68.7}$
$\mathbf{74.9}$
$\mathbf{79.0}$
$85.0$
Mixing Strategy. Although we have validated the effectiveness of our proposed FrustumMix in semi-supervised segmentation, we still compare FrustumMix with different popular outdoor mix strategies, including LaserMix [26], PolarMix [52]. As shown in Fig. 5 (b) and (c), the proposed FrustumMix significantly improves the performance and outperforms existing mixing strategies on both datasets.
Range-Interpolation Strategy. Range-Interpolation synthesizes possible points to reduce empty pixels in 2D representations. For the empty pixel in the range image, we initially use a $3\times 3$ window to pick its neighboring pixels and average valid pixel information as the interpolated point. Unfortunately, this results in a $0.4\%$ drop in performance. When we simply average the effective pixels, the generated points might fall out of the expected frustum region due to the empty surroundings. Based on this observation, we re-think the interpolation strategy. For an empty pixel, only when there are projection points on both the left and right or above and below, the calculated interpolation point can fall in the expected region. To this end, we propose two interpolation strategies along horizontal and vertical directions and improve the results by $0.4\%$ and $0.6\%$, respectively. It seems that points along horizontal directions share similar attributes as they line in the same laser beams.
@@
Backbones. To validate the versatility of our method, we select different range-view methods as our 2D backbone to extract frustum features, including SalsaNeXt [12], FIDNet [64], and CENet [8]. As shown in Fig. 5 (a), our proposed frustum-range representation can bring an improvement of $5.0\%$ to $6.8\%$ on different range-view methods.
5 Conclusion
@@
In this work, we presented a scalable Frustum-Range network (FRNet) for LiDAR segmentation, which can be seamlessly integrated into any range-view method. FRNet comprises three main components: the Frustum Feature Encoder Module to extract per-point features in low-level, the Frustum-Point Fusion Module to update per-point features hierarchically, and the Fusion Head Module to fuse different level features to predict more accurate labels. We also introduced two efficient augmentations, FurstumMix and Range-Interpolation, to improve the generalization of FRNet. Extensive experiments showed that our FRNet achieves competitive performance while maintaining high efficiency.
@@
[…]
Appendix A Additional Implementation Details
In this section, we provide additional details regarding the implementation and reproduction details of the proposed FRNet and Fast-FRNet.
A.1 Datasets
We adopt the following four LiDAR segmentation datasets to validate the effectiveness and superiority of the proposed approach.
  • SemanticKITTI [5] is a large-scale outdoor dataset which is based on KITTI odometry benchmark [14]. It consists of $22$ sequences, collected from different scenes in Karlsruhe, Germany by Velodyne HDL-64E. Commonly, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) is used for validation, and sequences $11$ to $21$ ($20,351$ scans) are used for testing. The original dataset is annotated with $28$ classes and only $19$ semantic classes are used for single-scan evaluation. The vertical field-of-view is $[-25^{\circ},3^{\circ}]$.
  • nuScenes [13] is a multimodal dataset that is widely used in autonomous driving. The LiDAR point clouds are collected around the streets in Boston and Singapore by Velodyne HDL-32E. It comprises $1,000$ driving scenes, in which $700$ scenes ($28,130$ scans) for training, $150$ scenes ($6,019$ scans) for validation, and $150$ scenes ($6,008$ scans) for testing. The dataset is annotated with $32$ classes and only $16$ semantic classes are used for evaluation. The vertical field-of-view is $[-30^{\circ},10^{\circ}]$.
  • ScribbleKITTI [49] is a scribble annotated dataset with line drawings, which shares the same LiDAR point clouds with SemanticKITTI [5]. It contains $189$ million labeled points, around $8.06\%$ of the total training points. This dataset is suitable for conducting weakly-supervised LiDAR segmentation, which is useful for reducing the annotation cost in existing LiDAR segmentation systems.
  • SemanticPOSS [37] is a more challenging dataset collected in Peking University by Pandora. The dataset includes $6$ sequences. Commonly, sequence $02$ ($500$ scans) is used for validation, and others ($2,488$ scans) are used for training. The dataset is annotated with $14$ classes and only $13$ semantic classes are used for evaluation. The vertical field-of-view is $[-16^{\circ},7^{\circ}]$.
A.2 Hyperparameters
For each point with coordinate $(x,y,z)$, its corresponding frustum region $(u,v)$ in the 2D representation with shape of $(H,W)$ is calculated as:
$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (11)
where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively. Points with the same $(u,v)$ are grouped into the same frustum region. The hyperparameter configuration of each module in our proposed FRNet framework is summarized as follows.
  • Frustum Feature Encoder. The input of frustum feature encoder for each point is $(x,y,z,i,d,x-x_{m},y-y_{m},z-z_{m})$, where $(x_{m},y_{m},z_{m})$ is the mean coordinate in the corresponding frustum region. The features are fed into four MLPs that map the channel from $8$ to $64$, $128$, $256$, and $256$, respectively, with batch normalization and ReLU function on the first three layers. Then the point features are pooled into frustum features, which are fed into an MLP that maps the channel from $256$ to $16$ with a ReLU function.
  • 2D Backbone. We adopt a variant of ResNet18 and ResNet34 used in CENet [8] as the backbone of Fast-FRNet and FRNet, respectively. The spatial shape of each stage is $[(H,W),(\frac{H}{2},\frac{W}{2}),(\frac{H}{4},\frac{W}{4}),(\frac{H}{8},\frac{W}{8})]$ and the output channel is $[128,128,128,128]$. The four features are interpolated to the shape of $(H,W)$ for feature concatenation, with two convolutional layers to reduce the channel dimension to $[256,128]$ progressively.
  • Frustum-Point Fusion Module. Point features are updated hierarchically at each stage of the 2D backbone. Accordingly, the channel dimension of point features at each stage is $[128,128,128,128]$. We also concatenate point features at each stage and reduce the channel dimension to $[256,128]$ with two MLPs progressively.
  • Optimization. For point-level supervision, we use the cross-entropy loss. For frustum-level supervision, we use cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42], and the weights for each loss are $1.0$, $1.5$, and $1.0$, respectively. The frustum-level loss is also used at each stage of the 2D backbone to supervise the middle frustum features. The weights for point-level loss and frustum-level loss are $1.0$ and $1.0$, respectively.
Algorithm 1 FrustumMix, NumPy-style
# lidar_list: the whole dataset.
# h: height of the 2D representation.
# w: width of the 2D representation.
# f_u: upward pitch angles of the field-of-view.
# f_d: downward pitch angles of the field-of-view.
# num_areas: List of split regions along inclination direction. We set [3, 4, 5, 6] here.
# sample: function to sample the lidar points and labels from the whole dataset.
# group: function to calculate frustum regions for points, refer to Eq. (1).
def FrustumMix(xa, ya, xb, yb):
    # group frustum regions
    ua, va = group(xa, h, w, f_u, f_d)
    ub, vb = group(xb, h, w, f_u, f_d)
    out_points = []
    out_labels = []
    if np.random.rand() > 0.5:
        # mix two scans along inclination direction
        n = np.random.choice(num_areas, size=1)[0]
        inc_list = np.linspace(0, h, n+1, dtype=np.int64)
        for i in range(n):
            start = inc_list[i]
            end = inc_list[i+1]
            if i % 2 == 0:
                idx = (va>=start) & (va<end)
                out_points.append(xa[idx])
                out_labels.append(ya[idx])
            else:
                idx = (vb>=start) & (vb<end)
                out_points.append(xb[idx])
                out_labels.append(yb[idx])
    else:
        # mix two scans along azimuth direction
        start = np.random.randint(0, w//2)
        end = start + w//2
        idxa = (ua<start) | (ua>=end)
        idxb = (ub>=start) & (ub<end)
        out_points = [xa[idxa], xb[idxb]]
        out_labels = [ya[idxa], yb[idxb]]
    out_points = np.concatenate(out_points)
    out_labels = np.concatenate(out_labels)
    return out_points, out_labels
# sample two sans
idx_a = np.random.randint(0, len(lidar_list))
idx_b = np.random.randint(0, len(lidar_list))
points_a, labels_a = sample(idx_a)
points_b, labels_b = sample(idx_b)
# mix the two scans
points, labels = FrustumMix(points_a, labels_a, points_b, labels_b)
Algorithm 2 Range-Interpolation, NumPy-style
# h: height of the range image.
# w: width of the range image.
# f_u: upward pitch angles of the field-of-view.
# f_d: downward pitch angles of the field-of-view.
# points: lidar points with shape [m, c].
# labels: semantic labels with shape: [m,].
# ignore_label: ignored label that is not used to calculate loss.
# laser_scan: function to convert LiDAR points to range image.
def RangeInterpolation(p, l):
    # convert to range images
    # r_m means the valid mask for each pixel. 0 is invalid, while 1 is valid.
    r_p, r_l, r_m = laser_scan(p, l, h, w, f_u, f_d)
    out_points = [p]
    out_labels = [l]
    for y in range(h):
        for x in range(w):
            if r_m[y, x]:
                continue
            if (x-1>=0) and (x+1<w):
                if r_m[y, x-1] and r_m[y, x+1]:
                    # both right and left pixels are valid
                    point = (r_p[y, x-1]+r_p[y, x+1]) / 2
                    if r_l[y, x-1] == r_l[y, x+1]:
                        label = r_l[y, x-1]
                    else:
                        label = ignore_label
                    out_points.append(point)
                    out_labels.append(label)
                    r_p[y, x] = point
                    r_l[y, x] = label
                    r_m[y, x] = 1
    out_points = np.concatenate(out_points)
    out_labels = np.concatenate(out_labels)
    return out_points, out_labels
points, labels = RangeInterpolation(points, labels)
A.3 Training Configurations
The common augmentation methods, including random dropping, random flipping, global rotation, global scale, global translation, and instance copy, are used in FRNet. Additionally, we also use the following proposed augmentation methods to further improve the performance.
  • FrustumMix. Based on the calculated frustum region $(u,v)$, the two scans are randomly mixed along inclination and azimuth directions. For inclination direction, the frustum regions are randomly split into several subsets sampled from $[3,4,5,6]$, while for azimuth direction, the frustum regions are split into $2$ subsets.
  • Range-Interpolation. Range-Interpolation aims to synthesize some possible points based on range images, which helps to fill empty pixels in the 2D representations. We first scan all the pixels in the range image. For each empty pixel, the interpolated points are calculated only when its right and left pixels both have projected points.
During the validation and testing phase, only the Range-Interpolation is kept to fill the empty pixels in 2D representations. To make the final results correspond to the original points, we set the first $N$ labels as the predicted labels as the last multiple points are interpolated points. Moreover, we apply the test time augmentation to improve the performance. Specifically, the input points are flipped along different directions, with randomly rotated, scaled, and jittered. The total number of voted point cloud scenes is $4$. For each scene, FRNet generates confidence scores for each class. We add the confidence scores of the $4$ scenes and then predict the final semantic labels.
Appendix B Additional Experimental Results
In this section, we provide the complete experimental results and qualitative visual examples among the four popular LiDAR segmentation benchmarks used in this work.
@@ B.1 Scalability and Robustness
B.2 Per-Class Results
B.2.1 SemanticKITTI
Tab. 5 shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [5]. We compare FRNet with different LiDAR representations, including point-view, range-view, voxel-view, and multi-view methods. Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods. Compared with SphereFormer [27] and UniSeg [32], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy. More results about efficiency and accuracy can be found in the main body. It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
B.2.2 nuScenes
Tab. 6 and Tab. 7 show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [13], respectively. The results demonstrate the great advantages of Fast-FRNet and FRNet. For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters. Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
B.2.3 ScribbleKITTI
To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [49], including RangeNet++ [35], SalsaNext [12], FIRNet [64], CENet [8], and RangeViT [4]. All the models are trained with their official settings. As shown in Tab. 8, FRNet achieves state-of-the-art performance. However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
B.2.4 SemanticPOSS
We also re-implement some popular voxel-view methods, including SPVCNN [47], Cylinder3D [65], and MinkNet [10], on SemanticPOSS [37]. All the experiments are conducted based on MMDetection3D framework [11]. Tab. 9 summarizes the class-wise IoU scores among these methods and some other popular range-view methods. Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods. For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.
B.3 Qualitative Results
In this section, we provide additional qualitative examples to further validate the effectiveness and superiority of the proposed FRNet framework.
B.3.1 Visual Comparisons
We supplement visual comparison examples of the proposed FRNet with prior works [4] and show the results in Fig. 9. Compared with range-view methods, FRNet can handle better for objects with complex structures, such as vegetation and terrain. In addition, the objects have much clearer boundaries.
B.3.2 Failure Cases
Although FRNet achieves great improvements among popular range-view methods, there are still some limitations that need to be considered. As shown in Fig. 8, FRNet tends to misclassify objects with similar structures to a large extent, i.e., bicyclist and motorcyclist, sidewalk, and other-ground. Spatial-temporal methods will help to alleviate this problem and yield better performance.
B.3.3 Video Demos
In addition to the figures shown in the main body and supplementary material, we also provide some video demos to support the superiority of FRNet. Specifically, we attached three video clips in the supplementary file, named demo1.mp4, demo2.mp4, and demo3.mp4. The first two demos include the error maps of the proposed FRNet and the prior art RangeViT [4]. The third demo contains the error maps and semantic predictions of FRNet. Each video demo shows LiDAR semantic segmentation sequences with hundreds of frames. All three demos validate again the effectiveness and superiority of FRNet in tackling the challenging LiDAR semantic segmentation task. We hope these video demos can provide a more comprehensive qualitative assessment of our proposed approach.
Table 5: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [5] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@ B.3.3 Video Demos
Table 6: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [13]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [9] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [35] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [61] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [38] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [12] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [63] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [4] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [65] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [31] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [54] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [24] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [27] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
@@
Table 7: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [13]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.

### v2

@@ IV-B Comparative Study
$62.7$ $72.3$ $72.8$ $82.7$
$64.3$ $73.5$ $73.6$ $83.4$
$66.2$ $73.9$ $76.4$ $84.6$
$67.0$ $74.1$ $77.2$ $84.8$
$67.6$ $74.2$ $77.7$ $\mathbf{85.2}$
$\mathbf{68.7}$ $\mathbf{74.9}$ $\mathbf{79.0}$ $85.0$
Quantitative Assessments. We compare the class-wise IoU scores of FRNet with those of the baseline method [58] in Fig. 5. Notably, FRNet demonstrates significant improvement across most semantic classes, particularly for dynamic classes with intricate structures, where it achieves mIoU gains ranging from $15\%$ to $24\%$. In Fig. 6, we present the prediction results on SemanticKITTI [28], compared with those of state-of-the-art range view methods, including FIDNet [55], CENet [58], and RangeViT [31]. It is evident that FRNet delivers more accurate predictions, particularly on large planar objects such as “vegetation”, “terrain”, and “sidewalk”, which are often challenging to distinguish accurately.
IV-C Ablation Study
In this section, we will discuss the efficacy of each design in our FRNet architecture. Unless otherwise specified, all experiments are conducted and reported on the val set of SemanticKITTI [28] and nuScenes [43], respectively.
Component Designs. Tab. IV summarizes the ablation results of each component in the FRNet architecture. Firstly, We directly apply the frustum feature encoder to CENet [58], incorporating KNN post-processing to predict semantic labels over the entire point cloud, yielding $62.7\%$ and $72.8\%$ mIoU on SemanticKITTI [28] and nuScenes [43], respectively. Subsequently, by integrating the point-frustum fusion to hierarchically update point features, the semantic results are predicted with the updated point features in an end-to-end manner, resulting in an improvement of $1.6\%$ and $0.8\%$ mIoU, respectively. To regularize the frustum features, the frustum-level supervision is introduced and it leads to significant gains of $1.9\%$ and $2.8\%$ mIoU, respectively. The fusion head module, which fuses multiple features at different levels for more accurate prediction, brings an additional $0.8\%$ mIoU gain on both datasets. Furthermore, the proposed RangeInterpolation, aimed at alleviating empty pixels in the 2D representation by reconstructing semantic surfaces via surrounding range information, provides a performance boost of approximately $0.5\%$ mIoU. Finally, adopting Test Time Augmentation during inference, following prior works, brings an improvement of $1.1\%$ and $1.3\%$ mIoU, respectively.
Figure 7: Ablation study on different FRNet configurations. Plots show the comparisons of: a) Different range-view backbones on the val set of SemanticKITTI [28]. b) Different mixing strategies on the val set of nuScenes [43]. c) Different mixing strategies on the val set of SemanticKITTI [28]. Color-filled boxes denote mAcc while empty boxes denote mIoU in (b) and (c), respectively.
Mixing Strategies. We further explore the effectiveness of the proposed FrustumMix. As shown in Tab. II, FrustumMix significantly improves performance on semi-supervised learning tasks and demonstrates generalizability to other modalities. To contextualize FrustumMix within common data augmentation techniques, we compare it with recent popular outdoor mixing strategies, including LaserMix [44] and PolarMix [45]. As illustrated in Fig. 7(b) and Fig. 7(c), FrustumMix consistently outperforms these prior strategies.
TABLE V: Ablation study on window settings in RangeInterpolation on the val set of SemanticKITTI [28]. Green pixel denotes the empty pixel needed to be interpolated. Violet pixels are the surrounding pixels used to generate the point. All mIoU scores are given in percentage (%).
Baseline $1\times 3$ $3\times 1$ $1\times 5$ $5\times 1$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$67.0$ $67.6_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.6)}}}}$ $67.4_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.4)}}}}$ $67.5_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.5)}}}}$ $67.4_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.4)}}}}$
cross $3\times 3$ cross $5\times 5$ $3\times 5$ $3\times 3$ $5\times 5$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$67.5_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.5)}}}}$ $67.4_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.4)}}}}$ $67.1_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.1)}}}}$ $66.8_{\mathbf{{{\color[rgb]{0.546875,0.33203125,0.609375}\definecolor[named]{pgfstrokecolor}{rgb}{0.546875,0.33203125,0.609375}(-0.2)}}}}$ $66.6_{\mathbf{{{\color[rgb]{0.546875,0.33203125,0.609375}\definecolor[named]{pgfstrokecolor}{rgb}{0.546875,0.33203125,0.609375}(-0.4)}}}}$
RangeInterpolation Window Settings. RangeInterpolation reconstructs semantic surfaces based on the range image, facilitating a compact 2D representation for coherent semantic learning. Tab. V showcases the performance of RangeInterpolation with different window settings. Notably, We observe consistent performance improvements when considering interpolation along the horizontal or vertical direction exclusively. Specifically, configurations employing horizontal windows, such as the $1\times 3$ and $1\times 5$ settings, demonstrate superior results compared to those focusing on the vertical direction ($3\times 1$ and $5\times 1$ windows). We conjecture that points aligned along the same laser beams often share similar attributes, leading to more accurate and coherent semantic surfaces. Conversely, larger window sizes, such as the $3\times 5$, $3\times 3$, and $5\times 5$ configurations, tend to exhibit non-uniform point distributions within the window, resulting in noisy points that fail to accurately represent the expected region and consequently leading to a drop in performance.
@@ IV-C Ablation Study
2D Backbones. To validate the versatility of our method, we employ various range-view methods as our 2D backbone for extracting frustum features, including SalsaNext [57], FIDNet [55], and CENet [58]. As depicted in Fig. 7(a), our proposed frustum-range representation consistently delivers performance gains ranging from $5.0\%$ to $6.8\%$ in mIoU across different range-view methods.
V Conclusion
@@
In this work, we presented FRNet, a Frustum-Range network designed for efficient and effective LiDAR segmentation, which can be seamlessly integrated into range-view approaches. FRNet comprises three key components: the Frustum Feature Encoder Module for extracting per-point features at lower levels, the Frustum-Point Fusion Module for hierarchical updating of per-point features, and the Fusion Head Module for fusing features from different levels to predict more accurate labels. Additionally, We introduced two efficient augmentation methods, FurstumMix and RangeInterpolation, to enhance the generalizability of range-view LiDAR segmentation. Extensive experiments across diverse driving perception datasets verified that FRNet achieves competitive performance while maintaining high efficiency. We hope this work can lay a solid foundation for future works targeted at real-time, in-vehicle 3D perception. Moving forward, we will further enhance the representation learning capability of FRNet and incorporate such a method into other 3D perception tasks, such as 3D object detection and semantic occupancy prediction.
@@
VI Additional Experimental Results
In this section, we provide the complete experimental results among the four popular LiDAR segmentation benchmarks used in this work.
@@ VI-A Scalability and Robustness
TABLE VI: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [28] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE VII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [43]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
Method (year) mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [88] [’21] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [30] [’19] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [33] [’20] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [81] [’22] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [57] [’20] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [18] [’22] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [31] [’23] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [16] [’21] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [20] [’20] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [21] [’21] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [32] [’23] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [68] [’23] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
@@
TABLE VIII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [43]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.